OpenStudy (anonymous):

Medal to best answer. Max and Maggie have to clean the house. It takes Max 12 hours to clean the house, while Maggie can complete the task in 4 hours. Their sister says that it will take 3 hours to complete if they work together. Explain each step in solving this equation, and determine if the sister is correct or not.

Vocaloid (vocaloid):

(1/12)x + (1/4)x = 1 solve for x

OpenStudy (anonymous):

48? @vocaloid

Vocaloid (vocaloid):

nope not even close lol start by multiplying both sides by 12

OpenStudy (anonymous):

haha i thought it sounded a little off. okay so does the 1/12 cancel out?

OpenStudy (anonymous):

@vocaloid

OpenStudy (anonymous):

wait aren't i supposed to find a common denominator?

OpenStudy (dariusx):

x + 3x = 12 4x = 12 x = 3

Vocaloid (vocaloid):

yeah, you could find a common denominator if you wanted it's just faster to multiply everything by 12 though the 1/12 cancels out what is (1/3)x * 12?

OpenStudy (anonymous):

12/3 = 4

Vocaloid (vocaloid):

what is (1/4)x times 12?

Vocaloid (vocaloid):

( meant to say 1/4, not 1/3)

OpenStudy (anonymous):

okay. it's 3 @vocaloid

Vocaloid (vocaloid):

don't forget the x

Vocaloid (vocaloid):

3x

Vocaloid (vocaloid):

so we have 3x + x = 12 what is x?

OpenStudy (anonymous):

x + 3x = 12 4x = 12 x = 3

Vocaloid (vocaloid):

good, so x = 3 since x represents the number of hours, her sister is correct

OpenStudy (anonymous):

Okay, so my answer sounds like this: Since x represents the number of hours, Max and Maggie's sister is correct. It will take both of them 3 hours to clean the house if they work together. First, make an equation. x + 3x = 12. Add the left portion together and you will get 4x = 12. Divide by four on both sides to isolate x and you will get x = 3. So, again it will take 3 hours for Maggie and Max to complete their work as their sister said.

Vocaloid (vocaloid):

not bad but I would start off with (1/12)x + (1/4)x = 1 as your equation